OpenStudy (anonymous):

y = 2x + 4 2x - y = 6 Solve the system of equations using substitution. A) (2,8) B) (2,-8) C) (10,24) D) no solution

OpenStudy (anonymous):

@Ahsome

OpenStudy (ahsome):

Oh, substitution. Have you done this before?

OpenStudy (anonymous):

i think its b or a

OpenStudy (ahsome):

How did you get that?

OpenStudy (anonymous):

hang on let me go over my work

OpenStudy (ahsome):

K

OpenStudy (anonymous):

i have no clue lol uhm d?

OpenStudy (ahsome):

lol Ok, lets see your two equations: \[y=2x+4\]\[2x-y=6\] Now, since you haven't done this, bear with me. In the first equation, we have \(y\) by itself on the left hand side. Now, I want to do the same thing for the second equation. It will be apparent soon why I want this. Lets just do this: \[2x-y=6\]\[-y=6-2x\]\[y=-1(6-2x)\]\[y=-6+2x\] Now, you can see, we have these two equations: \[y=2x+4\]\[y=-6+2x\] Can you see how they both equal \(y\)? Since they both equal \(y\), they must both be equal as well! THerefore, we can say: \[y=2x+4\text{ and }y=-6+2x\] \[2x+4=-6+2x\] Now, lets solve for \(x\) \[2x+4=-6+2x\]\[2x+10=2x\]\[10=2x-2x\]\[10=0\]But thats impossible, since \(10\ne 0\) Therefore, there is solution, since \(10\ne 0\), the answer is D :)

OpenStudy (anonymous):

ok thank you so much will you help me with 2 more?

OpenStudy (ahsome):

I will try, just got to eat lunch right now. After, I will, K?

OpenStudy (ahsome):

If you thinked I help, please consider pressing the Best Response button, Thanks :)

OpenStudy (anonymous):

lunch ok lol its 11 pm here and i will give you a medal

OpenStudy (ahsome):

Thanks man :)

OpenStudy (anonymous):

np